1)

The current in the conductor varies over time as

2)

The current in the conductor varies over time as

3)

The current in the conductor varies over time as

4)

The current in the conductor varies over time as


Help ASAP. will choose best answer quickly.
Answer #1

1). current = dq/dt

so Q= idt

so current passed during this time = area under the curve fro t=0 to t= 4s

so

charge transfereed=2*1+1/2*4*1+6*1.5+1/2*.5*4

= 14 Coulomb

b). i*4= 14

so i= 14/4

= 3.5 A


2). resistance of 9.77 cm of copper wire is R= rho*L/A

=1.68*10^-8*9.77*10^-2/[3.14*(1.35/2)^2*10^-4]

=1.1472*10^-5 ohm

so voltage difference= iR

=61*R

= 6.998*10^-4 V

=699.8 u V


3).dR = ?*dr/(2*?*r*L)


Integrate

?dR = ?/(2*?*L) * ?(1/r)*dr

R = ?/(2*?*L) * ln(r) [r = ra to rb)

R = ?/(2*?*L) * [ln(rb) - ln(ra)]

R = ?/(2*?*L) * ln(rb/ra)

R=3.9*10^5/(2*3.14*0.03)*ln(1.19/0.31)


R=2.78 Mohm


4).ner resistance is R= Ro(1-5*10^-4*(91-16)

Ro= 141

so R=135.72

so I= Vo/R

= 4.8/135.72

=0.0353688 A

= 35.368 mA


7). I= V/R

=16/39=0.41 A


8)n= i*t/1.6*10^-19

= 0.41*16*60/1.6*10^-19

= 2.46*10^21 electrons